OpenStudy (anonymous):

There is a sales tax of 18$ on an item that costs 222 before tax. The sales tax on a second item is 15 . How much does the second item cost before tax?

OpenStudy (kropot72):

First find the rate of sales tax as follows where t = sales tax rate as a decimal fraction: 18 = 222 * t t = 18/222 = 0.081 Let C = cost of second item. Then 15 = C * 0.081. Therefore C = 15/0.081 = $185.18

OpenStudy (anonymous):

icorrect thanks though

OpenStudy (anonymous):

incorrect*

OpenStudy (kropot72):

Doing the calculation without rounding the sales tax rate gives an answer of $185.00 as the cost of the second item before tax.